The bank contains 24 coins in nickels and quarters. The coins have a total value of $4.40. Find the number of nickels and the number of quarters in the banks
Answer by ewatrrr(24785)

Let x represent the number of Quarters (25 CENTS @)
25x + 5(24-x) = 440 CENTS makes sense
20x = 440 - 120
20x = 320
x = 16, number of Quarters. 8 Nickels 24-16
And...
25*16 + 5*8 = 440 CHECKS!